# Seeing through Imagination: Learning Scene Geometry via Implicit Spatial World Modeling

### v1

@@ Abstract
Figure 1: (a) Conceptual illustrations of our implicit spatial world modeling. (b) Example case of GeoGen dataset. Note that the point cloud scene is for illustrative purposes only, whereas raw video data is actually employed as input. (c) Performance boost of our MILO over different baselines including Video-3D LLM [zheng2025video], VGLLM [zheng2025learning], and RoboRefer [zhou2025roborefer].
@@ 1 Introduction
Spatial reasoning [yang2025thinking, zhang2025flatland, deng20253d] is the cognitive process of understanding and interpreting the 3D structure of the physical world, e.g., computing metric quantities and inferring spatial relationships among entities. It lies at the core of numerous vision and robotics tasks such as autonomous driving [shao2024lmdrive, sima2024drivelm], embodied navigation [zheng2024towards, lin2025navcot], and robotic manipulation [gao2024physically, nasiriany2024pivot, qu2025spatialvla]. Early attempts [fu2024scene, zhang2024chatscene, hong20233d] commonly utilize 3D point clouds as input and align them with the representation space of Multimodal Large Language Models (MLLMs). Nevertheless, due to the limited ability of current MLLMs to directly process 3D data and the high cost of acquiring high-quality 3D assets, recent efforts have shifted toward learning 3D priors from multi-view images or videos [zheng2025video, qi2025gpt4scene, zheng2025learning, fan2025vlm, chen2025sd].
@@
To enhance the spatial reasoning capabilities, most existing methods follow a verbal descriptive tuning paradigm [liao2025improved, chen2025sd, zhang2025flatland, zheng2025learning, fan2025vlm, huang2025mllms], i.e., they curate spatially-oriented datasets and instruct MLLMs to describe spatial properties (e.g., relative direction, object distance) in a single textual modality through supervised fine-tuning or reinforcement learning. Despite the progress, this descriptive tuning paradigm suffers from visual illiteracy, i.e., it relies solely on the textual symbolic supervision and is never exposed to how spatial transformations actually manifest in the visual domain. Consequently, MLLMs often fail to attend to the correct region of interest. In Figure 2, we visualize the attention scores of the last generated token with respect to all visual tokens. The baseline model, Video-3D LLM [zheng2025video], fails to properly attend to the region containing the target “wooden chair”, indicating that the current training scheme lacks cross-modal grounding between spatial semantics and visual perception. In contrast, for human spatial cognition, imagining and mentally simulating spatial structures is an intuitive process that grounds reasoning in perceptual experience rather than symbolic abstraction.
@@
Therefore, we propose an iMplicit spatIaL wOrld modeling (MILO) paradigm that complements current verbal descriptive tuning with visual generative tuning. As shown in Figure 1(a), a generator is integrated after MLLMs to yield visual feedback supervision under geometry-aware transformation instructions, such as viewpoint changes in Figure 1(b). In this way, we implicitly bridge symbolic spatial reasoning with perceptual grounding, enabling MLLMs to internalize how geometric transformations manifest in the visual domain. Notably, a recent work Ross3D [wang2025ross3d] also adopts a form of visual generative tuning; however, it is primarily confined to the tasks of masked visual prediction and BEV reconstruction. Moreover, it lacks explicit transformation instructions fed into the MLLMs, which prevents the models from perceiving and reasoning about the underlying geometric transformations they perform.
@@
Figure 2: Visualization of the attention scores of the last generated token with respect to all visual tokens. The baseline MLLM [zheng2025video] demonstrates visual illiteracy, i.e., failing to attend to the correct region of interest when generating responses. Refer to Table 5 for detailed quantitative analysis and supplementary material for more visualization cases.
@@
To inject 3D awareness into pre-trained MLLMs, recent works [zheng2025video, zhu2024llava] introduce 3D positional embeddings that encode patch-wise 3D coordinates in the world coordinate system and bind them to 2D semantic patch tokens [radford2021learning] to form “3D patches.” However, such absolute positional encodings inherently rely on a global video-specific coordinate system, making them sensitive to arbitrary world-coordinate choices and thus hindering generalization across datasets or camera setups. To this end, we propose a Relative Positional Encoding (RePE) scheme that captures relative camera-pose variations between adjacent frames based on intrinsic and extrinsic parameters. Unlike absolute encodings, our RePE is coordinate-system agnostic and explicitly models the geometric relationships among multiple views (e.g., how cameras are positioned and oriented with respect to one another).
@@
To facilitate the training of our MILO, we construct a Geometry-aware Generative (GeoGen) dataset, which models interaction outcomes under various geometric transformations (i.e., novel view synthesis and trajectory generation). As shown in Figure 1 (b), GeoGen consists of input visual frames (i.e., observations), the executed instructions (i.e., actions), and the resulting frames (i.e., outcomes). To ensure extensive coverage, we construct the GeoGen dataset with videos sourced from both scanned 3D scenes and the web, comprising 2,241 videos and 267,827 observation–action–observation triplets. By conducting our visual generative post-training on the GeoGen dataset followed by verbal descriptive tuning, we validate the consistent effectiveness of MILO across multiple baseline models, including Video-3D LLM [zheng2025video], VGLLM [zheng2025learning], and RoboRefer [zhou2025roborefer]. As shown in Figure 1 (c), when built upon Video-3D LLM, our MILO achieves a 3.2% absolute gain on the Acc@@0.25 metric of ScanRef [chen2020scanrefer] dataset.
@@ 2.1 Spatial Reasoning
Despite substantial progress in general scene understanding, MLLMs still struggle with 3D spatial reasoning [xu2025defining, zha2025enable, cai2025has, gholami2025spatial, yin2025spatial, ray2024sat, zhang2025open3dvqa, ma20243dsrbench, zhang2024sphere, lee2025perspective], which typically requires geometric and relationship reasoning in three-dimensional space. To mitigate this, preliminary approaches rely on the input point clouds [fu2024scene, hong20233d, zhang2024chatscene] or depth images [liu2025ssr, cai2024spatialbot, daxberger2025mm, chen2025sd, qi2025gpt4scene]. However, acquiring high-quality 3D assets is costly, and lifting 2D data into 3D [miao2025towards, chen2024spatialvlm] forms often involves complex pipelines, which severely impedes scalability and generalization. Therefore, recent studies have attempted to understand the 3D world directly from video data by curating large-scale, spatially oriented instruction datasets. To inject 3D priors, VGLLM [zheng2025learning] and VLM3R [fan2025vlm] augment the input with geometric embeddings extracted from pre-trained 3D foundation models [wang2025vggt, wang2025continuous]. Another line of work, including 3DRS [huang2025mllms] and ThinkWith3D [chen2025think], adopts a distillation strategy that leverages VGGT [wang2025vggt] as a teacher model to calibrate the visual embeddings of MLLMs. However, existing spatial instruction tuning remains confined to the verbal level, and how the visual hidden states of MLLMs contribute to the final response remains unclear. Therefore, we propose a visual generative tuning paradigm that supervises MLLMs with visual feedback, i.e., the visual states resulting from transformation instructions, which enables MLLMs to establish a human-like implicit spatial world modeling.
@@ 2.2 World Models
World Models [sutton1991dyna, ha2018recurrent, hafner2019dream, hafner2020mastering] are initially proposed to learn compact latent representations for visual dynamics prediction. Recently, with the remarkable generative capabilities of video diffusion models, emerging studies have begun exploring their potential as interactive world models capable of simulating complex and controllable environments [alonso2024diffusion, bruce2024genie, che2024gamegen, valevski2024diffusion, xiao2025worldmem]. By conditioning on the agent’s actions, these models can synthesize responsive and dynamic virtual worlds. On the other hand, such agent-environment interaction can also serve as an auxiliary prediction task for the agent’s policy [zhang2025agent, yu2025dyna]. Early Experience [zhang2025agent] leverages future states to enhance the agent by constructing internal representations of environmental dynamics, while Dyna-Think [yu2025dyna] integrates world model simulation directly into the agent’s reasoning process. For spatial reasoning tasks, WorldLM [zhang2025can] employs world model [blattmann2023stable] as dynamic state encoders to capture temporal variations. Although effective, it suffers from limited interpretability and incurs additional input embeddings. In contrast, our MILO framework actively enables geometry-aware interactions between MLLMs and the environment, thereby learning from the visual feedback of such interactions. The recent study Ross3D [wang2025ross3d] also adopts a generative fine-tuning paradigm, but it lacks explicit environment interaction since no concrete transformation instructions are provided to MLLMs. Moreover, it is confined to fixed masked visual prediction and BEV reconstruction tasks, resembling more a form of self-supervised and supervised representation learning.
@@ 3.1 Relative Positional Encoding
where we set $m=\lceil C/32\rceil$ to match the channel dimension of the semantic embeddings $\boldsymbol{E}_{i}^{\text{2D}}$ extract by CLIP [radford2021learning]. Then the final 3D-aware representation $\boldsymbol{E}_{i}$ for the $i$-the frame is computed as follows:
@@ 3.2 Implicit Spatial World Modeling
As illustrated in Figure 1, a video diffusion model $\mathcal{D}$ is used to reconstruct targets conditioned on the hidden visual embeddings of MLLM $\mathcal{P}$. Following the diffusion formulation, we adopt a denoising process that iteratively recovers clean latent tokens $z_{0}=\mathcal{F}(\boldsymbol{I})$ from noisy tokens $z_{t}$, where $\boldsymbol{I}$ denotes the target visual sequences and $\mathcal{F}(\cdot)$ is implemented as a continuous VAE provided by FLUX [labs2025flux1kontextflowmatching]. The diffusion model $\mathcal{D}$ is optimized to predict the noise $\epsilon$ given timestep $t$:
@@
Table 1: Experimental results on 3D scene understanding benchmarks. “Expert Models” train a separate model for each task, whereas “Generalist MLLMs” train a single model for all tasks. denotes reproduced results. Our MILO is built upon Video-3D LLM [zheng2025video].
@@
3D-VLP [jin2023context] 51.4 39.5 54.9 32.3 67.0 21.7
3D-VisTA [zhu20233d] 50.6 45.8 61.6 34.0 69.6 22.4 48.5
3DJCG [cai20223djcg] 49.6 37.3 26.6 49.5 31.0
@@
3D-LLM [hong20233d] 30.3 69.4 20.5 49.4
Chat-3D v2 [wang2023chat] 42.5 38.4 45.1 41.6 63.9 31.8 87.6 54.7
LL3DA [chen2024ll3da] 62.9 36.0 76.8
SceneLLM [fu2024scene] 80.0 27.2 53.6
LEO [huang2023embodied] 72.4 38.2 101.4 21.5 50.0
Grounded 3D-LLM [chen2024grounded] 47.9 44.1 45.2 40.6 70.6 35.5 72.7
PQ3D [zhu2024unifying] 57.0 51.2 50.1 80.3 36.0 47.1
ChatScene [zhang2024chatscene] 55.5 50.2 57.1 52.4 77.1 36.3 87.7 21.6 54.6
Ross3D [wang2025ross3d] 60.6 54.1 58.8 53.8 82.2 46.4 105.1 30.6 60.6
LLaVA-3D [zhu2024llava] 54.1 42.4 79.2 41.1 91.7 27.0 55.6
Inst3D-LLM [yu2025inst3d] 57.8 51.6 58.3 53.5 79.7 38.3 88.6 24.6
3D-LLaVA [deng20253d] 51.2 40.6 78.8 36.9 92.6 54.5
Video-3D LLM [zheng2025video] 58.1 51.7 58.0 52.7 83.8 41.3 102.1 30.1 58.6
@@ 3.3 GeoGen Dataset
  • Scanned 3D assets: We integrate public datasets from ScanNet [dai2017scannet] and ScanNet++ [yeshwanth2023scannetplus], which consist of RGB-D videos of indoor scenes annotated with 3D camera poses, reconstructed surface meshes, and instance-level semantic labels;
  • Internet videos: We leverage the RoomTour3D [han2025roomtour3d] dataset since it comprises geometric annotations including camera trajectories, relative depth maps, object tags, and bounding boxes.
@@
  • Novel view synthesis: As shown in Figure 6(a), given a reference frame and directional instructions, the model outputs corresponding video frames. To avoid abrupt scene transitions, we first partition each video into separate shots using PySceneDetect [PySceneDetect]. According to the camera pose information of the collected video data, we can easily derive relative camera positional relationships (e.g., forward/backward, left/right) between frames to construct the primary instructions.
  • Trajectory generation: As shown in Figure 6(b), given a complete input video, the model is required to predict the trajectory from a start point to an end point, necessitating a comprehensive understanding of the scene’s geometric structure. To achieve this, we propose a camera-calibrated graph based annotation method. Specifically, each video frame is treated as a node in the graph, and the edges between nodes are formed if their camera distance falls below a predefined threshold (i.e., 0.5m) and no obstructions are detected (verified via 3D point clouds and depth maps). By randomly selecting start and end nodes, we apply A search [hart1968formal] to determine the shortest path, with the sequence of traversed nodes constituting the trajectory. The primary transformation instruction is formulated as: “Please show me the path from {object in start frame} to {object in end frame}.”
Finally, the primary instruction–answer pairs and corresponding videos are fed to GPT-4o [gpt4o] to rephrase the instructions for greater diversity and richer scene details.
@@ 4.1 Experiment Settings
  • Video-3D LLM [zheng2025video]: This is one of the early attempts at 3D scene understanding. We replace the original 3D position encoding in Video-3D LLM with our proposed RePE and apply the visual generative post-training paradigm (i.e., Stage 1). Then, we follow the official settings of Video-3D LLM for fine-tuning (i.e., Stage 2). The experimental datasets include ScanRefer [chen2020scanrefer] and Multi3DRefer [zhang2023multi3drefer] for spatial referential understanding, Scan2Cap [chen2021scan2cap] for 3D scene captioning, and ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d] for spatial question answering.
  • VG-LLM [zheng2025learning]: It is a recent work that leverages VGGT [wang2025vggt] to augment MLLMs’ inputs. We firstly perform visual generative post-training (i.e., Stage 1). Following the original configuration, we use a sampled dataset from SPAR-7M [zhang2025flatland] and LLaVA-Hound split of LLaVA-Video-178K [zhang2024video] as the training dataset for the fine-tuning of stage 2. VSI-Bench [yang2025thinking] is used for evaluation.
  • RoboRefer [zhou2025roborefer]: This is designed for embodied spatial referring. Specifically, we firstly incorporate visual generative tuning into the depth alignment stage and then follow the official fine-tuning recipes. RefSpatial-Bench [zhou2025roborefer] is used for evaluation.
Notably, since VG-LLM [zheng2025learning] and RoboRefer [zhou2025roborefer] do not explicitly utilize 3D coordinates for input, we also omit the use of our RePE for a fair comparison.
Implementation Details. For visual generative tuning, the training step, batch size, and learning rate are set to 300, 16, and 1e-5 for the Video-3D LLM baseline [zheng2025video]; 500, 256, and 5e-6 for the VG-LLM baseline [zheng2025learning]; and 11534, 448, and 1e-3 for the RoboRefer baseline [zhou2025roborefer], respectively. During this stage, the MLLM backbone, the diffusion network, and the connector in-between are set to trainable while the visual encoder is frozen. All experiments are conducted on 8 NVIDIA A100 GPUs. For the verbal fine-tuning stage, we follow the original training recipes of the baselines [zheng2025video, zheng2025learning, zhou2025roborefer].
Table 2: Experimental results on VSI-Bench. denotes the reproduced results. Our MILO is build upon VG-LLM [zheng2025learning] with the same training dataset configurations.
@@
GPT-4o [gpt4o] - 34.0 46.2 5.3 43.8 38.2 37.0 41.3 31.5 28.5
Gemini-1.5-Flash [team2024gemini] - 42.1 49.8 30.8 53.5 54.4 37.7 41.0 31.5 37.8
Gemini-1.5-Pro [team2024gemini] - 45.4 56.2 30.9 64.1 43.6 51.3 46.3 36.0 34.6
InternVL2-8B [chen2024far] 8B 34.6 23.1 28.7 48.2 39.8 36.7 30.7 29.9 39.6
InternVL2-40B [chen2024far] 40B 36.0 34.9 26.9 46.5 31.8 42.1 32.2 34.0 39.6
LLaVA-NeXT-7B [li2024llavanext-strong] 7B 35.6 48.5 14.0 47.8 24.2 43.5 42.4 34.0 30.6
LLaVA-NeXT-72B [li2024llavanext-strong] 72B 40.9 48.9 22.8 57.4 35.3 42.4 36.7 35.0 48.6
LLaVA-OV-7B [li2024llava] 7B 32.4 47.7 20.2 47.4 12.3 42.5 35.2 29.4 24.4
LLaVA-OV-72B [li2024llava] 72B 40.2 43.5 23.9 57.6 37.5 42.5 39.9 32.5 44.6
Qwen2.5-VL-7B [bai2025qwen2] 7B 33.0 40.9 14.8 43.4 10.7 38.6 38.5 33.0 29.8
Qwen2.5-VL-72B [bai2025qwen2] 72B 37.0 25.1 29.3 54.5 38.8 38.2 37.0 34.0 28.9
@@
SPAR [zhang2025flatland] 8B 41.1 - - - - - - - -
Video-R1 [feng2025video] 7B 37.1 - - - - - - - -
vsGRPO-V [liao2025improved] 7B 40.7 59.9 29.6 50.8 48.3 35.4 35.6 34.0 31.5
SpaceR [ouyang2025spacer] 7B 45.6 - - - - - - - -
VG-LLM [zheng2025learning] 8B 59.5 69.9 51.4 66.5 64.2 63.5 79.6 45.4 28.3

### v2

@@ Abstract
Figure 1: (a) Conceptual illustrations of our implicit spatial world modeling. (b) Example case of GeoGen dataset. Note that the point cloud scene is for illustrative purposes only, whereas raw video data is actually employed as input. (c) Performance boost of our MILO over different baselines including Video-3D LLM [80], VGLLM [79], and RoboRefer [81].
@@ 1 Introduction
Spatial reasoning [64, 71, 21] is the cognitive process of understanding and interpreting the 3D structure of the physical world, e.g., computing metric quantities and inferring spatial relationships among entities. It lies at the core of numerous vision and robotics tasks such as autonomous driving [53, 54], embodied navigation [78, 41], and robotic manipulation [25, 46, 50]. Early attempts [24, 70, 32] commonly utilize 3D point clouds as input and align them with the representation space of Multimodal Large Language Models (MLLMs). Nevertheless, due to the limited ability of current MLLMs to directly process 3D data and the high cost of acquiring high-quality 3D assets, recent efforts have shifted toward learning 3D priors from multi-view images or videos [80, 49, 79, 22, 13].
@@
To enhance the spatial reasoning capabilities, most existing methods follow a verbal descriptive tuning paradigm [40, 13, 71, 79, 22, 34], i.e., they curate spatially-oriented datasets and instruct MLLMs to describe spatial properties (e.g., relative direction, object distance) in a single textual modality through supervised fine-tuning or reinforcement learning. Despite the progress, this descriptive tuning paradigm suffers from visual illiteracy, i.e., it relies solely on the textual symbolic supervision and is never exposed to how spatial transformations actually manifest in the visual domain. Consequently, MLLMs often fail to attend to the correct region of interest. In Figure 2, we visualize the attention scores of the last generated token with respect to all visual tokens. The baseline model, Video-3D LLM [80], fails to properly attend to the region containing the target “wooden chair”, indicating that the current training scheme lacks cross-modal grounding between spatial semantics and visual perception. In contrast, for human spatial cognition, imagining and mentally simulating spatial structures is an intuitive process that grounds reasoning in perceptual experience rather than symbolic abstraction.
@@
Therefore, we propose an iMplicit spatIaL wOrld modeling (MILO) paradigm that complements current verbal descriptive tuning with visual generative tuning. As shown in Figure 1(a), a generator is integrated after MLLMs to yield visual feedback supervision under geometry-aware transformation instructions, such as viewpoint changes in Figure 1(b). In this way, we implicitly bridge symbolic spatial reasoning with perceptual grounding, enabling MLLMs to internalize how geometric transformations manifest in the visual domain. Notably, a recent work Ross3D [58] also adopts a form of visual generative tuning; however, it is primarily confined to the tasks of masked visual prediction and BEV reconstruction. Moreover, it lacks explicit transformation instructions fed into the MLLMs, which prevents the models from perceiving and reasoning about the underlying geometric transformations they perform.
@@
Figure 2: Visualization of the attention scores of the last generated token with respect to all visual tokens. The baseline MLLM [80] demonstrates visual illiteracy, i.e., failing to attend to the correct region of interest when generating responses. Refer to Table 5 for detailed quantitative analysis and supplementary material for more visualization cases.
@@
To inject 3D awareness into pre-trained MLLMs, recent works [80, 82] introduce 3D positional embeddings that encode patch-wise 3D coordinates in the world coordinate system and bind them to 2D semantic patch tokens [51] to form “3D patches.” However, such absolute positional encodings inherently rely on a global video-specific coordinate system, making them sensitive to arbitrary world-coordinate choices and thus hindering generalization across datasets or camera setups. To this end, we propose a Relative Positional Encoding (RePE) scheme that captures relative camera-pose variations between adjacent frames based on intrinsic and extrinsic parameters. Unlike absolute encodings, our RePE is coordinate-system agnostic and explicitly models the geometric relationships among multiple views (e.g., how cameras are positioned and oriented with respect to one another).
@@
To facilitate the training of our MILO, we construct a Geometry-aware Generative (GeoGen) dataset, which models interaction outcomes under various geometric transformations (i.e., novel view synthesis and trajectory generation). As shown in Figure 1 (b), GeoGen consists of input visual frames (i.e., observations), the executed instructions (i.e., actions), and the resulting frames (i.e., outcomes). To ensure extensive coverage, we construct the GeoGen dataset with videos sourced from both scanned 3D scenes and the web, comprising 2,241 videos and 267,827 observation–action–observation triplets. By conducting our visual generative post-training on the GeoGen dataset followed by verbal descriptive tuning, we validate the consistent effectiveness of MILO across multiple baseline models, including Video-3D LLM [80], VGLLM [79], and RoboRefer [81]. As shown in Figure 1 (c), when built upon Video-3D LLM, our MILO achieves a 3.2% absolute gain on the Acc@@0.25 metric of ScanRef [12] dataset.
@@ 2.1 Spatial Reasoning
Despite substantial progress in general scene understanding, MLLMs still struggle with 3D spatial reasoning [63, 69, 8, 26, 66, 52, 75, 43, 74, 37], which typically requires geometric and relationship reasoning in three-dimensional space. To mitigate this, preliminary approaches rely on the input point clouds [24, 32, 70] or depth images [42, 7, 20, 13, 49]. However, acquiring high-quality 3D assets is costly, and lifting 2D data into 3D [45, 11] forms often involves complex pipelines, which severely impedes scalability and generalization. Therefore, recent studies have attempted to understand the 3D world directly from video data by curating large-scale, spatially oriented instruction datasets. To inject 3D priors, VGLLM [79] and VLM3R [22] augment the input with geometric embeddings extracted from pre-trained 3D foundation models [59, 60]. Another line of work, including 3DRS [34] and ThinkWith3D [18], adopts a distillation strategy that leverages VGGT [59] as a teacher model to calibrate the visual embeddings of MLLMs. However, existing spatial instruction tuning remains confined to the verbal level, and how the visual hidden states of MLLMs contribute to the final response remains unclear. Therefore, we propose a visual generative tuning paradigm that supervises MLLMs with visual feedback, i.e., the visual states resulting from transformation instructions, which enables MLLMs to establish a human-like implicit spatial world modeling.
@@ 2.2 World Models
World Models [55, 27, 28, 29] are initially proposed to learn compact latent representations for visual dynamics prediction. Recently, with the remarkable generative capabilities of video diffusion models, emerging studies have begun exploring their potential as interactive world models capable of simulating complex and controllable environments [1, 5, 10, 57, 62]. By conditioning on the agent’s actions, these models can synthesize responsive and dynamic virtual worlds. On the other hand, such agent-environment interaction can also serve as an auxiliary prediction task for the agent’s policy [72, 68]. Early Experience [72] leverages future states to enhance the agent by constructing internal representations of environmental dynamics, while Dyna-Think [68] integrates world model simulation directly into the agent’s reasoning process. For spatial reasoning tasks, WorldLM [73] employs world model [4] as dynamic state encoders to capture temporal variations. Although effective, it suffers from limited interpretability and incurs additional input embeddings. In contrast, our MILO framework actively enables geometry-aware interactions between MLLMs and the environment, thereby learning from the visual feedback of such interactions. The recent study Ross3D [58] also adopts a generative fine-tuning paradigm, but it lacks explicit environment interaction since no concrete transformation instructions are provided to MLLMs. Moreover, it is confined to fixed masked visual prediction and BEV reconstruction tasks, resembling more a form of self-supervised and supervised representation learning.
@@ 3.1 Relative Positional Encoding
where we set $m=\lceil C/32\rceil$ to match the channel dimension of the semantic embeddings $\boldsymbol{E}_{i}^{\text{2D}}$ extract by CLIP [51]. Then the final 3D-aware representation $\boldsymbol{E}_{i}$ for the $i$-the frame is computed as follows:
@@ 3.2 Implicit Spatial World Modeling
As illustrated in Figure 1, a video diffusion model $\mathcal{D}$ is used to reconstruct targets conditioned on the hidden visual embeddings of MLLM $\mathcal{P}$. Following the diffusion formulation, we adopt a denoising process that iteratively recovers clean latent tokens $z_{0}=\mathcal{F}(\boldsymbol{I})$ from noisy tokens $z_{t}$, where $\boldsymbol{I}$ denotes the target visual sequences and $\mathcal{F}(\cdot)$ is implemented as a continuous VAE provided by FLUX [36]. The diffusion model $\mathcal{D}$ is optimized to predict the noise $\epsilon$ given timestep $t$:
@@
Table 1: Experimental results on 3D scene understanding benchmarks. “Expert Models” train a separate model for each task, whereas “Generalist MLLMs” train a single model for all tasks. denotes reproduced results. Our MILO is built upon Video-3D LLM [80].
@@
3D-VLP [35] 51.4 39.5 54.9 32.3 67.0 21.7
3D-VisTA [83] 50.6 45.8 61.6 34.0 69.6 22.4 48.5
3DJCG [6] 49.6 37.3 26.6 49.5 31.0
@@
3D-LLM [32] 30.3 69.4 20.5 49.4
Chat-3D v2 [61] 42.5 38.4 45.1 41.6 63.9 31.8 87.6 54.7
LL3DA [14] 62.9 36.0 76.8
SceneLLM [24] 80.0 27.2 53.6
LEO [33] 72.4 38.2 101.4 21.5 50.0
Grounded 3D-LLM [15] 47.9 44.1 45.2 40.6 70.6 35.5 72.7
PQ3D [84] 57.0 51.2 50.1 80.3 36.0 47.1
ChatScene [70] 55.5 50.2 57.1 52.4 77.1 36.3 87.7 21.6 54.6
Ross3D [58] 60.6 54.1 58.8 53.8 82.2 46.4 105.1 30.6 60.6
LLaVA-3D [82] 54.1 42.4 79.2 41.1 91.7 27.0 55.6
Inst3D-LLM [67] 57.8 51.6 58.3 53.5 79.7 38.3 88.6 24.6
3D-LLaVA [21] 51.2 40.6 78.8 36.9 92.6 54.5
Video-3D LLM [80] 58.1 51.7 58.0 52.7 83.8 41.3 102.1 30.1 58.6
@@ 3.3 GeoGen Dataset
  • Scanned 3D assets: We integrate public datasets from ScanNet [19] and ScanNet++ [65], which consist of RGB-D videos of indoor scenes annotated with 3D camera poses, reconstructed surface meshes, and instance-level semantic labels;
  • Internet videos: We leverage the RoomTour3D [30] dataset since it comprises geometric annotations including camera trajectories, relative depth maps, object tags, and bounding boxes.
@@
  • Novel view synthesis: As shown in Figure 6(a), given a reference frame and directional instructions, the model outputs corresponding video frames. To avoid abrupt scene transitions, we first partition each video into separate shots using PySceneDetect [9]. According to the camera pose information of the collected video data, we can easily derive relative camera positional relationships (e.g., forward/backward, left/right) between frames to construct the primary instructions.
  • Trajectory generation: As shown in Figure 6(b), given a complete input video, the model is required to predict the trajectory from a start point to an end point, necessitating a comprehensive understanding of the scene’s geometric structure. To achieve this, we propose a camera-calibrated graph based annotation method. Specifically, each video frame is treated as a node in the graph, and the edges between nodes are formed if their camera distance falls below a predefined threshold (i.e., 0.5m) and no obstructions are detected (verified via 3D point clouds and depth maps). By randomly selecting start and end nodes, we apply A search [31] to determine the shortest path, with the sequence of traversed nodes constituting the trajectory. The primary transformation instruction is formulated as: “Please show me the path from {object in start frame} to {object in end frame}.”
Finally, the primary instruction–answer pairs and corresponding videos are fed to GPT-4o [47] to rephrase the instructions for greater diversity and richer scene details.
@@ 4.1 Experiment Settings
  • Video-3D LLM [80]: This is one of the early attempts at 3D scene understanding. We replace the original 3D position encoding in Video-3D LLM with our proposed RePE and apply the visual generative post-training paradigm (i.e., Stage 1). Then, we follow the official settings of Video-3D LLM for fine-tuning (i.e., Stage 2). The experimental datasets include ScanRefer [12] and Multi3DRefer [76] for spatial referential understanding, Scan2Cap [16] for 3D scene captioning, and ScanQA [2] and SQA3D [44] for spatial question answering.
  • VG-LLM [79]: It is a recent work that leverages VGGT [59] to augment MLLMs’ inputs. We firstly perform visual generative post-training (i.e., Stage 1). Following the original configuration, we use a sampled dataset from SPAR-7M [71] and LLaVA-Hound split of LLaVA-Video-178K [77] as the training dataset for the fine-tuning of stage 2. VSI-Bench [64] is used for evaluation.
  • RoboRefer [81]: This is designed for embodied spatial referring. Specifically, we firstly incorporate visual generative tuning into the depth alignment stage and then follow the official fine-tuning recipes. RefSpatial-Bench [81] is used for evaluation.
Notably, since VG-LLM [79] and RoboRefer [81] do not explicitly utilize 3D coordinates for input, we also omit the use of our RePE for a fair comparison.
Implementation Details. For visual generative tuning, the training step, batch size, and learning rate are set to 300, 16, and 1e-5 for the Video-3D LLM baseline [80]; 500, 256, and 5e-6 for the VG-LLM baseline [79]; and 11534, 448, and 1e-3 for the RoboRefer baseline [81], respectively. During this stage, the MLLM backbone, the diffusion network, and the connector in-between are set to trainable while the visual encoder is frozen. All experiments are conducted on 8 NVIDIA A100 GPUs. For the verbal fine-tuning stage, we follow the original training recipes of the baselines [80, 79, 81].
Table 2: Experimental results on VSI-Bench. denotes the reproduced results. Our MILO is build upon VG-LLM [79] with the same training dataset configurations.
@@
GPT-4o [47] - 34.0 46.2 5.3 43.8 38.2 37.0 41.3 31.5 28.5
Gemini-1.5-Flash [56] - 42.1 49.8 30.8 53.5 54.4 37.7 41.0 31.5 37.8
Gemini-1.5-Pro [56] - 45.4 56.2 30.9 64.1 43.6 51.3 46.3 36.0 34.6
InternVL2-8B [17] 8B 34.6 23.1 28.7 48.2 39.8 36.7 30.7 29.9 39.6
InternVL2-40B [17] 40B 36.0 34.9 26.9 46.5 31.8 42.1 32.2 34.0 39.6
LLaVA-NeXT-7B [38] 7B 35.6 48.5 14.0 47.8 24.2 43.5 42.4 34.0 30.6
LLaVA-NeXT-72B [38] 72B 40.9 48.9 22.8 57.4 35.3 42.4 36.7 35.0 48.6
LLaVA-OV-7B [39] 7B 32.4 47.7 20.2 47.4 12.3 42.5 35.2 29.4 24.4
LLaVA-OV-72B [39] 72B 40.2 43.5 23.9 57.6 37.5 42.5 39.9 32.5 44.6
Qwen2.5-VL-7B [3] 7B 33.0 40.9 14.8 43.4 10.7 38.6 38.5 33.0 29.8
Qwen2.5-VL-72B [3] 72B 37.0 25.1 29.3 54.5 38.8 38.2 37.0 34.0 28.9
@@
SPAR [71] 8B 41.1 - - - - - - - -
Video-R1 [23] 7B 37.1 - - - - - - - -
vsGRPO-V [40] 7B 40.7 59.9 29.6 50.8 48.3 35.4 35.6 34.0 31.5
SpaceR [48] 7B 45.6 - - - - - - - -
VG-LLM [79] 8B 59.5 69.9 51.4 66.5 64.2 63.5 79.6 45.4 28.3